Multiply out and simplify the following:
9(-10t-5)=
What is it?

Answer by ewatrrr(24785)

 
Hi,
Multiply out and simplify the following:
'Distributing the 9 '
9(-10t-5)= -90t - 45 = -45(2t+1)